OpenStudy (anonymous):

The temperature at which the tungsten filament of a 12V and 36W lamp operates is 1730 degrees celsius. If the temperature coefficient of resistance of tungsten is 6 x 10^ -3 /K, find the resistance of the lamp at a room temperature of 20 degrees celsius.

OpenStudy (anonymous):

See the formula here --> http://www.allaboutcircuits.com/vol_1/chpt_12/6.html Given, T= 1730 degree C = 2003 K We know that P=V*I Therefore I=P/V = 36/12 = 3 A R= V/ I = 12/3= 4 ohm Therefore 4= (R_ref)[1+ (6 x 10^ -3)(2003-293)] R_ref = 4/(1+(0.006)(1710)) = 0.35 ohms I hope this is what you are looking for

OpenStudy (anonymous):

thanks man